# Trial and Trust: Addressing Byzantine Attacks with Comprehensive Defense Strategy

### v1

@@ 5  Experiments
Adam 0.956$\pm$0.017 0.811$\pm$0.016 0.262$\pm$0.023 0.041$\pm$0.019 0.348$\pm$0.011 0.126$\pm$0.016 0.197$\pm$0.027 0.036$\pm$0.015 0.125$\pm$0.011 0.123$\pm$0.020
FLTrust 0.952$\pm$0.020 0.800$\pm$0.019 0.952$\pm$0.016 0.753$\pm$0.011 0.617$\pm$0.020 0.174$\pm$0.019 0.061$\pm$0.017 0.125$\pm$0.015 0.017$\pm$0.013 0.123$\pm$0.018
Recess 0.949$\pm$0.016 0.783$\pm$0.019 0.366$\pm$0.019 0.128$\pm$0.020 0.593$\pm$0.020 0.163$\pm$0.020 0.493$\pm$0.019 0.112$\pm$0.015 0.450$\pm$0.014 0.127$\pm$0.018
Zeno 0.921$\pm$0.012 0.787$\pm$0.014 0.014$\pm$0.017 0.110$\pm$0.015 0.163$\pm$0.010 0.089$\pm$0.014 0.102$\pm$0.012 0.066$\pm$0.018 0.010$\pm$0.009 0.091$\pm$0.011
CC 0.949$\pm$0.020 0.772$\pm$0.019 0.285$\pm$0.018 0.114$\pm$0.020 0.580$\pm$0.019 0.155$\pm$0.020 0.084$\pm$0.019 0.014$\pm$0.020 0.530$\pm$0.018 0.154$\pm$0.020
CC+fbm 0.954$\pm$0.016 0.808$\pm$0.020 0.840$\pm$0.019 0.716$\pm$0.014 0.562$\pm$0.011 0.151$\pm$0.020 0.027$\pm$0.018 0.123$\pm$0.015 0.876$\pm$0.017 0.594$\pm$0.013
CC+bucketing 0.947$\pm$0.013 0.790$\pm$0.018 0.829$\pm$0.011 0.708$\pm$0.020 0.570$\pm$0.012 0.164$\pm$0.018 0.035$\pm$0.020 0.118$\pm$0.012 0.870$\pm$0.019 0.587$\pm$0.014
Safeguard 0.957$\pm$0.020 0.821$\pm$0.019 0.107$\pm$0.012 0.123$\pm$0.020 0.258$\pm$0.011 0.124$\pm$0.019 0.951$\pm$0.018 0.082$\pm$0.020 0.010$\pm$0.009 0.123$\pm$0.012
VR Marina 0.010$\pm$0.014 0.120$\pm$0.010 0.027$\pm$0.018 0.123$\pm$0.020 0.176$\pm$0.012 0.103$\pm$0.013 0.127$\pm$0.013 0.079$\pm$0.019 0.012$\pm$0.010 0.108$\pm$0.013
BANT 0.953$\pm$0.017 0.830$\pm$0.020 0.956$\pm$0.016 0.777$\pm$0.020 0.948$\pm$0.018 0.809$\pm$0.020 0.946$\pm$0.020 0.676$\pm$0.015 0.947$\pm$0.018 0.770$\pm$0.020
AutoBANT 0.953$\pm$0.019 0.781$\pm$0.020 0.790$\pm$0.020 0.276$\pm$0.020 0.946$\pm$0.019 0.748$\pm$0.018 0.942$\pm$0.020 0.690$\pm$0.020 0.892$\pm$0.016 0.585$\pm$0.020
SimBANT 0.956$\pm$0.020 0.790$\pm$0.018 0.949$\pm$0.020 0.774$\pm$0.020 0.945$\pm$0.020 0.712$\pm$0.018 0.955$\pm$0.020 0.783$\pm$0.018 0.946$\pm$0.019 0.705$\pm$0.020
Method Without Attack ALIE (40%) Adam 41.14 $\pm$ 4.20 62.14 $\pm$ 4.83 FLTrust 62.79 $\pm$ 14.53 84.57 $\pm$ 15.17 Recess 86.55 $\pm$ 5.08 98.55 $\pm$ 3.24 Zeno 50.26 $\pm$ 4.35 67.23 $\pm$ 8.42 CC 66.90 $\pm$ 7.65 94.72 $\pm$ 6.17 Safeguard 53.29 $\pm$ 5.11 71.23 $\pm$ 4.67 BANT (ours) 46.17 $\pm$ 3.17 70.21 $\pm$ 5.13 AutoBANT (ours) 62.05 $\pm$ 7.67 83.62 $\pm$ 1.27 SimBANT (ours) 55.47 $\pm$ 6.34 72.49 $\pm$ 7.65
Table 2: Time per communication round for ResNet1d18 on ECG (AFIB)
Figure 1: Test accuracy for ResNet18 on CIFAR-10.
Figure 2: Test accuracy for ResNet18 on Dirichlet.
We define the number of Byzantine clients as a percentage of the total number of clients, and specify it in the attack name. We train BANT and AutoBANT in the scaled version (see Section 4.3) with the Adam preconditioner. We also include SimBANT, Adam, and the existing methods: Zeno (Xie et al., 2019), Recess (Yan et al., 2024), Centered Clip (Karimireddy et al., 2021b), Safeguard (Allen-Zhu et al., 2020), VR Marina (Gorbunov et al., 2021a) and FLTrust (Cao et al., 2020). For Centered Clip, we also added techniques Fixing by Mixing (Allouah et al., 2023) and Bucketing (Karimireddy et al., 2021a). The methods were trained on the CIFAR-10 and ECG datasets for 200 and 150 rounds, respectively. More technical details are presented in Appendix A.1.
The accuracy and loss curves with classification metrics for all considered attacks on the CIFAR-10 test dataset are illustrated in Figures 1, 3 and Table 3 (Appendix A.2), respectively. Figure 1 (see captions (g) and (h)) also shows the accuracy plots of the proposed methods under the Dirichlet distribution with $\alpha=0.5$ for the ALIE and Random Gradients attacks. More details are presented in the Appendix A.3. To assess model performance on the ECG data, we use the G-mean (the square root of sensitivity multiplied by specificity) and the f1-score metrics. Table 1 summarizes the results of the methods for the AFIB disease classification. Detailed results for all considered abnormalities across multiple metrics are presented in Tables 7-10 in Appendix A.2.
Unlike previously established techniques, our methods exhibit robustness against all Byzantine attacks. To further stress test the proposed methods, we consider the most strong Byzantine attacks under statistically heterogeneous setups modeled by the Dirichlet distribution. We also address time per communication round for the methods in Table 2 for the ECG setup. Note that our methods have training times per round comparable to baselines. Round times for the CIFAR-10 setup are presented in Table 4 in Appendix A.2.
We note that AutoBANT performs slightly worse compared to BANT and SimBANT under Random Gradients and ALIE attacks. This occurs due to solving an auxiliary subproblem (Line 10 in Algorithm 2) using Mirror Descent with KL-divergence. According to its properties, the algorithm assigns small but non-zero weights to Byzantines, contributing to unstable convergence, while BANT and SimBANT lack this drawback.
Recess and FLTrust leverage the concept of trust scores but rely on the majority of honest devices. As a result, it leads to a significant decrease in final quality under the majority of Byzantines in Random Gradients and IPM setups, as well as with ALIE attack that simulates a malicious majority. Similar behavior is observed for the CC and Safeguard methods, which also suffer from sensitivity to parameter tuning. Fixing by Mixing and Bucketing increase Label Flipping and ALIE metrics for ECG setup, but do not provide reliable convergence for all cases. Zeno exploits the trial function approach, but it relies on the number of Byzantine clients. We address the choice of this hyperparameter in Table 5 (Appendix A.2). In contrast, BANT-based methods demonstrate resilience to extreme attacks while enabling adaptive selection of trusted clients.
Acknowledgements
A significant part of this work was carried out while G. Molodtsov and D. Medyakov were visiting research assistants at Mohamed bin Zayed University of Artificial Intelligence (MBZUAI).
Appendix
Appendix A Additional Experiments
A.1  Technical details
Our implementation is developed in Python 3.10. We simulate a distributed system on a single server. The server is equipped with an AMD EPYC 7513 32-Core Processor running at 2.6 GHz and Nvidia A100 SXM4 40GB. This configuration is used for the experiments described in Section 5.
To ensure a fair comparison, we maintain consistent hyperparameters across all methods. The batch size is set to 32 for CIFAR-10 and 64 for the ECG dataset, with a local client learning rate of 0.003 and adam preconditioner. All clients perform local computations for 1 epoch. For CIFAR-10, cross-entropy was used as the local client loss function, while for ECG is was its binary version. To address the problem of imbalance of positive and negative examples in the ECG, the minority class was reweighted to the majority in the corresponding loss function. For the ECG classification task, we train the models on 10-second 12-lead ECG records, with all records resampled to a frequency of 500 Hz. Additionally, we train the models exclusively on patients older than 18 years of age.
We also select specific parameters for the implemented methods. For the Safeguard method, we use window sizes of 1 and 6 for two different accumulation settings, with the threshold chosen automatically, as described in the original paper. We adapt the CC method to a local computation case and set the clipping coefficient $\tau=0.1$ and the SGD momentum $\beta=0.9$ for CIFAR-10, as well as $\tau=1$ and $\beta=0.5$ for the ECG case. For both setups we fixed the number of clipping iterations to $l=1$. For the Fixing-by-Mixing technique, we set the number of Byzantine clients $f$ to be less than half of all clients as suggested in the article: for the CIFAR-10 $f=4$, while for the ECG $f=2$. For the Bucketing technique we apply 2-bucketing strategy with global learning rate $\eta=0.9$. For the Recess method, we set the decrease score equal to $0.1$. For the Zeno method, we set the regularization weight $\rho=0.0005$ as a default value in the paper and the trim parameter $b=2$ for the ECG setup and $b=3$ for CIFAR-10. We address the choice of hyperparameter $b$ in Table 5 in Appendix A.2, as it is critical in real-world scenarios. For the BANT method, we set the momentum parameter $\beta=0.5$. The AutoBANT method uses the number of optimization epochs equal to 5 and $\gamma=1$. For SimBANT, we set the softmax temperature parameter for the model logits to 0.05.
$sim_{\text{ECG}}(x,y)=1-|x-y|,$ (2)
$sim_{\text{CIFAR}}(x,y)=\frac{x\cdot y}{\|x\|\|y\|},$ (3)
where $y$ is one-hot encode targets.
A.2  CIFAR-10 Experiments
Table 3: ResNet18 on CIFAR-10. Accuracy under various attacks.
Algorithm Without Attack ALIE (40%) Adam 17.76 ± 1.76 30.74 ± 1.87 FLTrust 57.64 ± 7.55 72.75 ± 1.82 Recess 65.88 ± 16.02 76.99 ± 6.27 Zeno 38.13 ± 2.26 54.88 ± 4.71 CC 22.95 ± 2.30 43.34 ± 6.28 CC+fbm 23.44 ± 2.81 47.16 ± 7.95 CC+bucketing 20.18 ± 2.12 43.46 ± 6.19 Safeguard 43.26 ± 3.14 58.29 ± 2.31 VR Marina 55.84 ± 9.14 73.91 ± 6.89 BANT 29.31 ± 2.56 55.89 ± 6.29 AutoBANT 49.25 ± 5.77 67.27 ± 3.49 SimBANT 33.40 ± 5.00 60.11 ± 8.70
Table 4: Time (in seconds) per communication round for ResNet18 on CIFAR-10 without attack and under ALIE attack.
Table 4 reports the average time per communication round, including standard deviation for various federated learning algorithms using ResNet18 on CIFAR-10 in two settings: Without attack and ALIE. We see that Recess took the longest time, since it requires double local calculations of the client in one round of communication. AutoBANT requires a little more time compared to baselines, except for the Recess and FLTrust methods. We attribute this to the solution of the auxiliary task Algorithm 4.3, Line 10), which is slightly longer than in the ECG setup and has a greater impact on the execution time. However, it is not expected that the size of $\hat{f}$ increases significantly as the system scales. Thus, this contribution is negligible in real-world scenarios, as demonstrated by the ECG experiment in Table 2. In addition, these computations are performed on the central node, which has more computational resources in the federated learning paradigm. All of this reflects the applicability of the proposed methods in real-world setups.
A.3  Dirichlet Distributions
To examine the robustness of the proposed methods under a heterogeneous setup, we perform Byzantine attacks on CIFAR-10 with Dirichlet distribution of client data. So $\mathbf{p}_{k}\sim\text{Dir}(\alpha\cdot\mathbf{1}_{n})$ represents distribution of samples with label $k=\overline{1,K}$ over $n$ clients, i.e. $\mathbf{p}_{k}\in S_{n}(1)$ where $S_{n}(1)$ is the standard $n$ unit simplex. The hyperparameter $\alpha$ is called the concentration parameter, and it controls the degree of heterogeneity of the client data distribution. We address two heterogeneity regimes: medium with $\alpha=1$ and strong with $\alpha=0.5$. Since the goal of the experiment is to test the robustness of the proposed methods, we consider the 3 strongest attacks in our setup: Random Gradient (50%), ALIE (40%), and IPM (70%). Figure 4, 5 illustrates the performance of the BANT-based methods.
A.4  ECG Experiments
Now, let us move on to the choice of the trim hyperparameter $b$ in the Zeno method. This parameter cuts off the number of clients that will not be used in the aggregation of the global model in the training round. The ranking of clients is based on $Score_{\gamma,\rho}$, which exploits the trial function approach. In the paper, it is assumed that $b\geqslant q$, where $q$ is the number of Byzantines participating in the training. Thus, the method provides protection against an arbitrary number of malicious participants and can be stable in critical attack setups. However, in real-world scenarios, we do not have a priori information about the number of Byzantines. As a result, Zeno does not show comparable quality metrics in the ECG setup for $b=2$ compared to the BANT-like methods, which is reflected in Table 1.
Algorithm
Without Attack
G-mean f1-score
Label Flipping (60%)
G-mean f1-score
Random Gradients (60%)
G-mean f1-score
IPM (80 %)
G-mean f1-score
ALIE (40 %)
G-mean f1-score
Zeno ($b$ = 3) 0.953$\pm$0.018 0.806$\pm$0.017 0.950$\pm$0.020 0.753$\pm$0.019 0.954$\pm$0.020 0.770$\pm$0.017 0.945$\pm$0.017 0.730$\pm$0.020 0.946$\pm$0.015 0.717$\pm$0.016
Zeno 0.921$\pm$0.012 0.787$\pm$0.014 0.014$\pm$0.017 0.110$\pm$0.015 0.163$\pm$0.010 0.089$\pm$0.014 0.102$\pm$0.012 0.066$\pm$0.018 0.010$\pm$0.009 0.091$\pm$0.011
Table 5: ResNet1d18 on ECG (AFIB). G-mean and f1-score for Zeno with different hyperparameters under attacks.
We address the issue of choosing this hyperparameter in the ECG setup in Table 5. As can be seen from the results, a suitably selected trim coefficient value of $b=3$, which cuts off $60\%$ of Byzantine clients, shows comparable results with the BANT-like methods in all considered attack scenarios.
@@ A.4  ECG Experiments
Adam 0.304$\pm$0.015 0.116$\pm$0.018 0.952$\pm$0.014 0.738$\pm$0.011
FLTrust 0.586$\pm$0.018 0.179$\pm$0.015 0.011$\pm$0.014 0.123$\pm$0.013
Recess 0.359$\pm$0.018 0.115$\pm$0.014 0.933$\pm$0.017 0.611$\pm$0.018
Zeno 0.017$\pm$0.016 0.130$\pm$0.018 0.010$\pm$0.018 0.140$\pm$0.020
CC 0.479$\pm$0.020 0.124$\pm$0.017 0.945$\pm$0.020 0.710$\pm$0.016
CC+fbm 0.155$\pm$0.016 0.124$\pm$0.017 0.948$\pm$0.018 0.695$\pm$0.020
CC+bucketing 0.137$\pm$0.017 0.119$\pm$0.010 0.944$\pm$0.016 0.689$\pm$0.013
Safeguard 0.084$\pm$0.016 0.014$\pm$0.018 0.109$\pm$0.010 0.123$\pm$0.016
VR Marina 0.096$\pm$0.017 0.078$\pm$0.019 0.098$\pm$0.011 0.110$\pm$0.014
BANT 0.943$\pm$0.019 0.792$\pm$0.019 0.949$\pm$0.020 0.704$\pm$0.017
AutoBANT 0.737$\pm$0.020 0.243$\pm$0.019 0.948$\pm$0.018 0.695$\pm$0.015
SimBANT 0.951$\pm$0.020 0.760$\pm$0.020 0.965$\pm$0.017 0.753$\pm$0.020
Here, we will present all the obtained metric results for the four heart abnormalities we investigated: Atrial FIBrillation (AFIB), First-degree AV block (1AVB), Premature Ventricular Complex (PVC), and Complete Left Bundle Branch Block (CLBBB). In Table 6, we present results for AFIB pathology under 2 additional attacks scenarios: Sign Flipping $(60\%)$ and IPM $(60\%)$. In addition, we compare all methods for the attacks described in 5 section in Tables 7 - 10. To obtain confidence intervals in Tables 1, 5, 6, each run was repeated $5$ times. We did not estimate the variance of metrics for the 1AVB, PVC and CLBBB pathologies in Tables 8-10 due to computational challenges.
@@
Without Attack Adam 0.940 0.974 0.956 0.811
@@
Label Flip (60%) Adam 0.089 0.774 0.262 0.041
@@
Sign Flip (60%) Adam 0.096 0.961 0.304 0.116
@@
Random Gradients (60%) Adam 0.136 0.893 0.348 0.126
@@
Without Attack Adam 0.896 0.871 0.884 0.335
@@
Label Flip (60%) Adam 0.210 0.002 0.022 0.069
@@
Sign Flip (60%) Adam 0.597 0.685 0.640 0.119
@@
Random Gradients (60%) Adam 0.220 0.005 0.036 0.074
@@
Without Attack Adam 0.977 0.974 0.975 0.790
@@
Label Flip (60%) Adam 0.639 0.710 0.673 0.180
@@
Sign Flip (60%) Adam 0.639 0.710 0.673 0.180
@@
Random Gradients (60%) Adam 0.220 0.005 0.035 0.120
@@
Without Attack Adam 0.979 0.957 0.968 0.508
@@
Label Flip (60%) Adam 0.220 0.065 0.120 0.115
@@
Sign Flip (60%) Adam 1.000 0.001 0.027 0.044
@@
Random Gradients (60%) Adam 1.000 0.007 0.085 0.044
@@
Table 10: ResNet1d18 on ECG (CLBBB).
@@ Appendix B Notation
The following sections will be dedicated to the theoretical proof of all aspects discussed in the main body of the paper. A complete table of contents can be found on page 13.
@@
$N$ Number of samples in the trial function $\gamma$ Learning rate (stepsize) in optimization $\hat{f}$ Trial loss function on the server $\mathbb{I}_{[a>0]}$ Indicator function taking value 1 if $a>0$, otherwise 0
@@
$n=n(t)$ Number of workers in the distributed system $\mu$ Strong convexity constant (2.2(a)) $f_{i}$ Local objective function on $i$-th worker $\langle a,b\rangle_{\hat{P}}$ Weighted inner product with $\hat{P}$
$\mathcal{G}=\mathcal{G}(t)$ Set of honest workers $\beta$ Momentum parameter for weights $\omega_{i}^{t}$ $f_{1}$ Local objective function on the server $\alpha$ Lower bound on the preconditioner (F.1)
$G=G(t)$ Number of honest workers $\delta$ Approximation error in $\arg\min$ finding (2) $g_{i}$ Stochastic gradient of worker $i$ $\Gamma$ Upper bound on the preconditioner (F.1)
$\mathcal{B}=\mathcal{B}(t)$ Set of Byzantine workers $\delta_{1}$ $(\delta_{1},\delta_{2})-$heterogeneity parameter (2.4) $g_{i}^{t}$ Gradient from worker $i$ at iteration $t$$\cdot\|_{\infty}$ $\max\limits_{1\leqslant i\leqslant d}|\cdot|_{i}$
$B=B(t)$ Number of Byzantine workers $\delta_{2}$ $(\delta_{1},\delta_{2})-$heterogeneity parameter (2.4) $x^{*}$ Optimal solution of the objective function $\Delta_{d}^{1}$ $d$-dimensional simplex constraint on weights
@@
$T$ Total number of iterations in training $l$ Local round length $\mathcal{KL}(p\|q)$ Kullback-Leibler divergence between distributions $p$ and $q$ $[\cdot]_{0}$ Non-negative projection: $\max\{0,\cdot\}$
@@
[…]
Table 11: Notation Reference.

### v5

@@ A.4 ECG Experiments
[…]
Metrics for all abnormalities.
Here, we present all the obtained metric results for the four heart abnormalities we investigated: Atrial FIBrillation (AFIB), First-degree AV block (1AVB), Premature Ventricular Complex (PVC), and Complete Left Bundle Branch Block (CLBBB). In Table 8, we show results for AFIB pathology under 2 additional attacks scenarios: Sign Flipping $(60\%)$ and IPM $(60\%)$. In addition, we compare all methods for the attacks described in 6 section in Tables 9 - 12. To obtain confidence intervals in Tables 1, 7, 8, each run was repeated $5$ times. We do not report the variance of metrics for the 1AVB, PVC and CLBBB pathologies in Tables 10-12 due to overabundance and computational factor.
@@
[…]
Table 9: ResNet1d18 on ECG (AFIB).
@@
Without Attack
Adam 0.940 0.974 0.956 0.811
@@
Label Flip (60%)
Adam 0.089 0.774 0.262 0.041
@@
Sign Flip (60%)
Adam 0.096 0.961 0.304 0.116
@@
Random Gradients (60%)
Adam 0.136 0.893 0.348 0.126
@@
Without Attack
Adam 0.896 0.871 0.884 0.335
@@
Label Flip (60%)
Adam 0.210 0.002 0.022 0.069
@@
Sign Flip (60%)
Adam 0.597 0.685 0.640 0.119
@@
Random Gradients (60%)
Adam 0.220 0.005 0.036 0.074
@@
Without Attack
Adam 0.977 0.974 0.975 0.790
@@
Label Flip (60%)
Adam 0.639 0.710 0.673 0.180
@@
Sign Flip (60%)
Adam 0.639 0.710 0.673 0.180
@@
Random Gradients (60%)
Adam 0.220 0.005 0.035 0.120
@@
Without Attack
Adam 0.979 0.957 0.968 0.508
@@
Label Flip (60%)
Adam 0.220 0.065 0.120 0.115
@@
Sign Flip (60%)
Adam 1.000 0.001 0.027 0.044
@@
Random Gradients (60%)
Adam 1.000 0.007 0.085 0.044
@@
A.5 Learning-to-rank Experiments
In the main body, we introduced our approach and presented a subset of results for the Learning-to-Rank (LTR) task under Byzantine settings. Here, we expand on the experimental setup, covering all baseline models and proposed methods across the full spectrum of adversarial scenarios considered.
Problem Formulation.
The LTR task is defined over a set of queries $\mathcal{Q}$, where each query $q\in\mathcal{Q}$ is associated with a set of documents $D_{q}$. For each document $d_{i}\in D_{q}$, a feature vector $x_{i}$ and a relevance label $y_{i}\in\{0,\dots,r-1\}$ are provided. The objective is to learn a scoring function $f(x;\theta)$ such that, for any query $q$, the induced ordering of scores $s_{i}=f(x_{i};\theta)$ approximates the ideal relevance ordering.
We use the Normalized Discounted Cumulative Gain at cutoff $k$ (NDCG@k) to assess ranking quality. First, we define the Discounted Cumulative Gain (DCG@k) for a ranking $\pi$ (a permutation of documents based on predicted scores) as:
$\mathrm{DCG}@k=\sum_{i=1}^{k}\frac{2^{y_{\pi(i)}}-1}{\log_{2}(i+1)},$
where $y_{\pi(i)}$ is the relevance label of the document ranked at position $i$. This formulation assigns higher weight to highly relevant documents that appear earlier in the ranking, with a logarithmic discount applied to lower positions.
Figure 9: Test NDCG@5 for Transformer on the Learning-to-Rank task.
The Ideal DCG (IDCG@k) is the maximum possible DCG@k obtained by sorting documents in descending order of their true relevance labels.
The final evaluation metric, Normalized DCG (NDCG@k), is computed as:
$\mathrm{NDCG}@k=\frac{\mathrm{DCG}@k}{\mathrm{IDCG}@k}.$
This normalization bounds the metric between 0 and 1, where 1 indicates a perfect ranking.
Experimental Scope.
We evaluate models on the WEB30K dataset [DBLP:journals/corr/QinL13], a standard benchmark for LTR consisting of 30,000 queries with graded relevance labels. While the main section highlights our method’s applicability to LTR, here we provide complete validation across:
  • Baseline defenses, including Zeno, Recess, Centered Clip, Safeguard and FlTrust.
  • Our proposed methods: Bant, AutoBant, and SimBant.
  • All adversarial scenarios introduced in the main paper: Label Flipping, Sign Flipping, Random Gradients, IPM, and ALIE.
For consistency with Zeno’s assumptions, we set the threshold for Byzantine tolerance to $|\mathcal{B}|=0.5n$, where $n$ is the number of clients as it is unknown a priori. While Zeno might perform well when the fraction of adversarial clients is close to $50\%$, it suffers significant degradation (or divergence) when this assumption is violated. This sensitivity underscores the need for defenses that do not rely on tight prior knowledge of Byzantine ratios.
@@ Appendix B Notation
The following sections will be dedicated to the theoretical proof of all aspects discussed in the main body of the paper.
@@
Table 13: Notation Reference.
NN Number of samples in the trial function $\gamma$ Learning rate (stepsize) in optimization $\hat{f}$ Trial loss function on the server $\mathbb{I}_{[a>0]}$ Indicator function taking value 1 if $a>0$, otherwise 0
@@
$n=n(t)$ Number of workers in the distributed system $\mu$ Strong convexity constant (3.2(a)) $f_{i}$ Local objective function on $i$-th worker $\langle a,b\rangle_{\hat{P}}$ Weighted inner product with $\hat{P}$
$\mathcal{G}(t)$ Set of honest workers $\beta$ Momentum parameter for weights $\omega_{i}^{t}$ $f_{1}$ Local objective function on the server $\alpha$ Lower bound on the preconditioner (F.1)
$G(t)=|\mathcal{G}(t)|$ Number of honest workers $\delta$ Approximation error in $\arg\min$ finding (2) $g_{i}$ Stochastic gradient of worker $i$ $\Gamma$ Upper bound on the preconditioner (F.1)
$\mathcal{B}(t)$ Set of Byzantine workers $\delta_{1}$ $(\delta_{1},\delta_{2})$-heterogeneity parameter (3.4) $g_{i}^{t}$ Gradient from worker $i$ at iteration $t$ $\|\cdot\|_{\infty}$ $\max\limits_{1\leqslant i\leqslant d}|\cdot|_{i}$
$B(t)=|\mathcal{B}(t)|$ Number of Byzantine workers $\delta_{2}$ $(\delta_{1},\delta_{2})$-heterogeneity parameter (3.4) $x^{*}$ Optimal solution of the objective function $\Delta_{d}^{1}$ $d$-dimensional simplex constraint on weights
@@
$T$ Total number of iterations in training $l$ Local round length $\mathcal{KL}(p\|q)$ Kullback-Leibler divergence between distributions $p$ and $q$ $\cdot]_{0}$ Non-negative projection: $\max\{0,\cdot\}$
@@
$\mathcal{W}=\mathcal{W}(t)$ Number of workers $G$ $\underset{t\leqslant T}{\min}~G(t)$ $S$ Bound on $\mathcal{X}$ in Lemma C.1 $\|\cdot\|$ If not specified other, $\|\cdot\|_{2}=\sqrt{\langle\cdot,\cdot\rangle}$